OpenStudy (death47):

Divide. Express your answer in lowest terms. 1 1 -- / -- 5 3

TheSmartOne (thesmartone):

\(\LARGE \frac{a}{b} \div \frac{c}{d} = \frac{a}{b} \times \frac{d}{c} = \boxed{\frac{ad}{bc}}\)

OpenStudy (anonymous):

when dividing fractions, you reciprocal the term on the right, and multiply. Ex: 1/3 / 1/7 = 1/3 X 7/1 = 7/3

OpenStudy (anonymous):

flip the second one and multiply

OpenStudy (death47):

So that answer would be 3/5

OpenStudy (anonymous):

i have heard elementary school math teachers call this KFC Keep, Flip (the second one), Change ( the division to a multiplication)

OpenStudy (death47):

yea I get it, I just want someone to make sure I'm right :)

TheSmartOne (thesmartone):

3/5 is correct

OpenStudy (anonymous):

@Death47 u r right

OpenStudy (death47):

ok thank you guys :D

TheSmartOne (thesmartone):

Anytime! :)